# Scalable LinUCB: Low-Rank Design Matrix Updates for Recommenders with Large Action Spaces

### v1

@@ 7.1 Scalability with context size
Figure 1: Performance comparison across different context sizes on Magazine Subscriptions dataset .
7.2 Scalability with the number of arms
This experiment evaluates how the algorithms scale with the increasing number of arms.
We fix the optimal context dimensions for each dataset (Magazine, Health, Beauty, and Musical Instruments) using outer-product features selected by scree plot (see Section˜A.1.1) and choose the optimal PSI rank by cross-validaton (see Section˜A.1.1)). The number of arms is gradually increased from a small subset to the full action space. At each stage, for each state we select items with the highest observation frequency. Hyperparameters are tuned via cross-validation for each dataset and model (see Section˜A.1.2).
We compare PSI with two baselines, LinUCB and LinUCB Classic, measuring hit rate, computation time, and memory usage. As shown in Table˜3 and Figure˜27, both baseline methods fail to run at larger scales due to memory constraints, while PSI remains stable and achieves comparable perfomance with significantly lower time and memory costs.
@@ 7.2 Scalability with the number of arms
7.3 Recommendation quality vs. rank
To analyze how the rank parameter affects recommendation quality, we conducted an experiment in which we varied the rank of the PSI method while keeping the context dimensions and number of arms fixed for each dataset. We measured the quality of recommendations after the warm-up phase. At each rank configuration, we measured the hit rate and compared it against the LinUCB baseline. Each month we continued to train the model further and then measured its performance by calculation the relative hit rate to provide a demonstration of the improvement in the quality of our algorithm.
The experiment was conducted using the optimal hyperparameters (see  Section˜A.1.1 and Section˜A.1.2).
The results show that, starting from a relatively low rank, our method performs well across all datasets. This indicates that our approach is capable of maintaining high-quality recommendations without requiring a full-rank matrix.
Refer to caption Refer to caption
(a) MovieLens 1M (b) Amazon All Beauty
Refer to caption Refer to caption
(c) Amazon Health (d) Magazine Subscriptions
Figure 3: Quality comparison by months: PSI vs LinUCB
7.4 CBSCFD vs PSI
@@
This experiment compares the performance of the CBSCFD and PSI-LinUCB algorithms across multiple datasets. In the original CBSCFD method, a shared design matrix $A$ is used for all arms, which does not allow personalization. We extend this approach by training a separate matrix $A_{a}$ for each arm $a$, as proposed in li2010contextual .
CBSCFD supports only rank-1 updates, where a single context arrives at each time step, while PSI-LinUCB generalizes to support batch-context updates, where multiple contexts can be accumulated and updated at once.
Table 2 presents the hit rate performance across four datasets (MovieLens 1M, Amazon All Beauty, Amazon Health, and Magazine Subscriptions) for both algorithms with varying parameter configurations. PSI-LinUCB consistently outperforms CBSCFD.
@@
Figure 4 shows the relationship between the parameter $m$ and hit rate performance for CBSCFD across the four real-world datasets. The results show significant variation in the optimal $m$ values across different datasets, with unpredictable behavior. Even small changes in $m$ lead to varying results.
Figure 5 provides a corresponding analysis for PSI-LinUCB, demonstrating a consistent improvement in performance as the rank increases. The algorithm shows predictable behavior: the hit rate increases monotonically with rank. This trend holds across all datasets, including sparse matrices. The relationship between rank and performance remains stable, independent of dataset characteristics.
In the CBSCFD algorithm, the parameter $\alpha$ (which corresponds to $\varepsilon$ in our notations) is dynamically updated at each round: the initial value is obtained by accumulating the sum of lost singular values that are subtracted during matrix compression. This works as adaptive regularization that compensates for the lost information and maintains the positive definiteness of the approximation matrices. In our PSI-LinUCB algorithm, the regularization parameter $\varepsilon$ is fixed to 1; making it dynamic could improve stability by keeping the design matrix better conditioned, as classical LinUCB is sensitive to this parameter (the algorithms have the same optimal hyperparametres).
Table 2: Hit rate performance across datasets and algorithms
Dataset M1 M2 M3 M4 M5 M6 M7 M8 Avg
MovieLens 1M
PSI-LinUCB 0.065 0.102 0.075 0.054 0.041 0.046 0.048 0.046 0.060
CBSCFD 0.065 0.102 0.075 0.054 0.041 0.045 0.045 0.040 0.058
Amazon All Beauty
PSI-LinUCB 0.012 0.009 0.022 0.015 0.014 0.003 0.011 0.008 0.012
CBSCFD 0.008 0.010 0.022 0.017 0.019 0.002 0.011 0.011 0.012
Amazon Health
PSI-LinUCB 0.044 0.024 0.018 0.018 0.037 0.01 0.017 0.012 0.022
CBSCFD 0.044 0.024 0.014 0.018 0.037 0.01 0.01 0.007 0.020
Magazine Subscriptions
PSI-LinUCB 0.603 0.528 0.578 0.504 0.570 0.605 0.565 0.286 0.530
CBSCFD 0.595 0.538 0.586 0.496 0.561 0.613 0.548 0.257 0.524
Refer to caption Refer to caption
(a) MovieLens 1M (b) Amazon All Beauty
Refer to caption Refer to caption
(c) Amazon Health (d) Magazine Subscriptions
@@
Refer to caption Refer to caption
(a) MovieLens 1M (b) Amazon All Beauty
Refer to caption Refer to caption
(c) Amazon Health (d) Magazine Subscriptions
Figure 5: Average Hit rate for different rank values for PSI-LinUCB.
@@ 8 Conclusion
We presented Scalable-LinUCB, a computationally efficient variant of LinUCB that maintains a dynamically low-rank representation of the inverse Cholesky-style symmetric factor of the regularized design matrix. The method allows to control rank growth of the accumulated context matrix with a projector–splitting step. Moreover, our method has an average complexity of updates of order $\mathcal{O}(dr)$ per candidate. One of the potential future research directions involves substituting the projector–splitting integrator with other dynamical low-rank updates.
@@ Acknowledgements
Appendix A Experiments
A.1 Hyperparameter selection
A.1.1 Context dimensionality $r^{\prime}$.
@@ A.1 Hyperparameter selection
The optimal values of $r^{\prime}$ for context size were selected using the scree plot, which displays the singular values in descending order, indicating the optimal number of components that capture the most informative structure of the data.
Dataset Optimal $r^{\prime}$ Context size PSI rank $r$
Movielens 1M 13 169 32
Magazine Subscriptions 41 1681 64
Health & Personal Care 37 1369 128
All Beauty 51 2601 128
Musical Instruments 70 4900 256
A.1.2 Regularization $\alpha$.
@@
We used a grid search to find the optimal exploration parameter $\alpha$. First, a broad grid
$\alpha\in\{-10,-8,-6,-4,-2,0,2,4,6,8,10\}\,\,(\text{step size}\,2)$ (41)
was evaluated to assess the overall impact of different values of $\alpha$ on the model’s performance. Based on the results from this initial search, a more refined range was selected:
$\alpha=\{-0.1,\,-1,\,-0.5,\,-0.3,\,-0.2,\,-0.05,\,-0.03,\,-0.01,\,0.1\}.$ (42)
For each candidate $\alpha$, models were trained on the training split and selected by the best validation metric. Final performance is reported on the held-out test set using the selected $\alpha$.
The optimal exploration parameter $\alpha$ for the LinUCB coincided with the optimal $\alpha$ for PSI-LinUCB -LinUCB algorithm.
A.1.3 Rank grids for PSI-LinUCB -LinUCB and $m$ for CBSCBD
Model capacity for both the PSI-LinUCB -LinUCB rank and the CBSCBD parameter $m$ was explored over a following grid :
$\text{Ranks}=\bigl\{\,b\cdot 2^{\,i}\;\big|\;i\in\{1,2,3,4,5\},\;b\in\{2,3\}\,\bigr\}.$ (43)
Selection followed the same validation protocol as for $\alpha$.
Table 3: Performance comparison for the increasing number of arms with rank $r$
Dataset Context Items Algorithm Hit Rate Train Time (s) Train Mem (MB) Eval Time (s) Eval Mem (MB)
Magazine d=41 (1681) r=64 1000 PSI-LinUCB -LinUCB 0.5957 52.33 1593.5 13.95 1726.0
LinUCB 0.5957 1367.94 43173.8 40.09 43519.1
LinUCB Classic 0.5957 446.70 21761.9 39.85 21934.4
2000 PSI-LinUCB 0.5533 57.99 1748.9 28.81 1901.7
LinUCB 0.5533 1743.39 84730.6 90.61 86585.7
LinUCB Classic 0.5533 485.78 42597.7 89.63 43525.8
3000 PSI-LinUCB 0.5482 59.77 1794.4 44.21 1971.3
LinUCB 0.5482 1848.22 124381.9 139.33 129707.1
LinUCB Classic 0.5482 503.49 62436.4 139.69 65113.4
Health d=37 (1369) r=128 1000 PSI-LinUCB 0.0296 271.36 3189.5 13.56 3718.5
LinUCB 0.0296 1719.74 28496.6 20.48 29239.3
LinUCB Classic 0.0296 844.69 14514.1 21.05 14885.6
10000 PSI-LinUCB 0.0189 512.89 6683.4 311.16 7726.5
LinUCB failed - - - -
LinUCB Classic 0.0189 1627.97 133827.8 713.53 143940.4
60000 PSI-LinUCB 0.0167 639.88 9662.3 2576.1 11146.8
LinUCB failed - - - -
LinUCB Classic failed - - - -
Beauty d=51 (2601) r=128 1000 PSI-LinUCB 0.0401 447.50 6290.4 16.47 7313.7
LinUCB 0.0401 10575.42 100768.9 57.98 103970.1
LinUCB Classic 0.0401 4636.82 50695.9 56.71 52297.3
70000 PSI-LinUCB 0.0142 1265.85 24840.5 5507.90 28724.0
LinUCB failed - - - -
LinUCB Classic failed - - - -
110000 PSI-LinUCB 0.0122 1328.77 27205.7 9189.7 31562.6
LinUCB failed - - - -
LinUCB Classic failed - - - -
Musical Instruments d=70 (4900) r=256 1000 PSI-LinUCB 0.2574 3858.76 32758.2 811.49 34460.6
LinUCB failed - - - -
LinUCB Classic 0.2574 75289.6 183968.5 4089.5 185250.7
80000 PSI-LinUCB 0.0606 13652.76 161977.2 131747.78 186052.1
LinUCB failed - - - -
LinUCB Classic failed - - - -
200000 PSI-LinUCB 0.0458 14535.0 192015.1 189541.5 201018.1
LinUCB failed - - - -
LinUCB Classic failed - - - -
Figure 6: Algorithm scaling with number of arms on Health & Personal Care dataset.
@@ A.1.3 Rank grids for PSI-LinUCB -LinUCB and $m$ for CBSCBD
Figure 7: Algorithm scaling with number of arms on Magazine Subscriptions dataset.
Appendix B Proofs
B.1 Proof for Section˜4.2

### v2

@@ 5 Scalability results
Figure 1: Performance comparison across different context sizes on Amazon Health dataset.
Table 3: Hit Rate comparison across context dimensions $d$
$d$ LinUCB LinUCB Woodbury PSI-LinUCB
900 0.026 0.026 0.026
2500 0.034 0.034 0.034
4900 - - 0.044
16900 - - 0.050
Increasing number of arms. We fix the optimal context dimensions and PSI rank for each dataset (Health, Beauty, and Magazine Subscriptions), then gradually increase the number of arms from a small subset to the full action space, selecting items with the highest observation frequency at each stage. As shown in Figure˜2, exact implementations of LinUCB become grossly memory-inefficient as number of arms increases, while PSI-LinUCB remains computationally tractable. Additional results on time and memory consumption during the warm-up phase on the Beauty dataset (Figure 10(c)), as well as further scalability experiments are reported in the Appendix,  Appendix˜C.
@@
LinUCB quality. We analyze how the rank parameter affects recommendation quality by varying the PSI rank while keeping context dimensions and number of arms fixed. At each configuration, we measure hit rate of our algorithm, popular and random baselines relative to the one of LinUCB. Figure˜3 shows that PSI-LinUCB achieves recommendation quality comparable to classical LinUCB starting from relatively low ranks, without requiring a full-rank matrix. Additional results are provided in Appendix, see Figure˜8.
(a) Amazon Health
(b) Amazon All Beauty
Figure 3: Quality comparison by months: PSI-LinUCB vs LinUCB on Amazon datasets.
6 Comparison with scalable baselines
6.1 Online setting
Following the experimental setup of kuzborskij2019efficient and chen2020efficient, we evaluate our method on an online classification problem turned into contextual bandit setting. To maintain compatibility with the mentioned works, in this subsection we switch to the version of LinUCB with shared parametrization, see e.g. chen2020efficient with a single parameter vector and design matrix across all arms.
Given a dataset with $K$ classes, we mark of them as a target. In each round, the environment randomly draws one sample from each class, forming a context set. The learner selects one sample and receives a reward of $1$ if the selected sample belongs to the target cluster, and 0 otherwise. Note that in this binary reward setting, cumulative number of mistakes corresponds to the cumulative regret. We conduct experiments on two benchmark datasets: MNIST LeCun1998Gradient and CIFAR-10 Krizhevsky2009Learning. For fair comparison we tuned the optimal rank/sketch size for all algorithms via cross-validation by minimizing mistakes over $T$ rounds. The dataset statistics are summarized in the Appendix, Table˜6, the detailed description of the experimental setup and hyperparameter selection is provided in Section˜6.3. Since this experiment focuses on online learning, we use the rank-1 version of PSI-LinUCB (see Algorithm˜6 in Appendix), which is consistent with the update scheme employed by CBSCFD chen2020efficient, CBRAP yu2017cbrap, and DBSL wen2024matrix.
The results, presented in Figure 4, demonstrate that our PSI-LinUCB algorithm achieves competitive performance with CBSCFD on MNIST while being significantly faster and outperforms all other baseline methods in both quality and computational efficiency. Results on CIFAR dataset are presented in Appendix, Appendix˜C, Figure 11).
Figure 4: Online classification results on the MNIST dataset.
6.2 Approximation quality of the inverse matrix
Figure 5 shows the dependence of the relative approximation error of the inverse design matrix $A_{t}^{-1}$ on final iteration for different ranks (sketch sizes for CBSCFD). The approximation error of PSI-LinUCB and CBSCFD decreases as the rank increases. However, PSI-LinUCB consistently achieves lower error than CBSCFD for the same rank size. The corresponding experiments on CIFAR (see Figure 16) and real-world datasets (see Figures 16(a), 16(b)) are presented in Appendix,  Appendix˜C.
Figure 5: Approximation error of $A_{t}^{-1}$ on MNIST
6.3 Performance on real-world datasets
@@
This experiment compares PSI-LinUCB with CBSCFD, CBRAP and DBSLinUCB across multiple datasets. In the original methods, a shared design matrix $A_{t}$ is used for all arms; we extend this by training a separate matrix $A_{t,a}$ for each arm $a$, as proposed in li2010contextual. Additionally, CBSCFD and CBRAP support only rank-1 updates, while PSI-LinUCB generalizes to batch updates. The experiment follows the setup described in Section˜4.1, for full details see Section˜C.1 in Appendix.
Table 4 presents quality metrics across three datasets (Amazon Magazine Subscriptions, Amazon All Beauty, Amazon Health), and Table 5 reports training time, per-user prediction time and memory usage. CBRAP achieves quality comparable to PSI-LinUCB and CBSCFD but requires a larger sketch size $m$, resulting in higher memory and time costs. This explains inherently large training and evaluation time of CBRAP on the Beauty dataset. DBSLinUCB introduces additional tuning complexity, requiring joint search over the initial block size $l_{0}$ (which grows exponentially) and the error parameter $\varepsilon$, with interdependent effects. For example, DBSLinUCB fails on the Health dataset with all checked hyperparameters, see Table˜4. CBSCFD shows comparable quality to PSI-LinUCB, yet we point out its main drawback as instability of the quality with respect to sketch size $m$.
@@
Figure˜6 and Figure˜7 show the relationship between the parameter $m$ and hit rate for CBSCFD and PSI-LinUCB, respectively. The results demonstrate significant variation in the optimal $m$ values across different real-world datasets, with unpredictable behavior: even small changes in $m$ lead to different results. At the same, for PSI-LinUCB the dependence of quality on approximation rank $r$ is monotone as the rank increases, which is desirable for tuning this (arguably, most important) hyperparameter in practice. We provide additional experiments on other datasets in the Appendix, Appendix˜C (see Figure 14).
Table 4: Quality metrics on different datasets
Dataset Algorithm Hit@10 NDCG@10 MRR@10 Cov.
Magazine PSI-LinUCB 0.531 0.430 0.398 0.038
CBSCFD 0.526 0.449 0.425 0.031
CBRAP 0.528 0.420 0.386 0.027
DBSLinUCB 0.523 0.438 0.411 0.028
Health PSI-LinUCB 0.022 0.017 0.016 0.0016
CBSCFD 0.020 0.015 0.014 0.0012
CBRAP 0.020 0.017 0.016 0.0012
DBSLinUCB 0.001 0.001 0.001 0.0006
Beauty PSI-LinUCB 0.013 0.011 0.010 0.0076
CBSCFD 0.012 0.009 0.009 0.0011
CBRAP 0.012 0.011 0.010 0.0011
DBSLinUCB 0.013 0.011 0.010 0.0047
Table 5: Computational efficiency on different datasets
Dataset Algorithm Train (s) Eval (s) Train (GB) Eval (GB)
Magazine PSI-LinUCB 10.3 0.145 1.4 1.6
CBSCFD 22.9 0.163 1.5 1.5
CBRAP 134.9 9.224 1.5 1.5
DBSLinUCB 24.6 0.102 1.5 1.8
Health PSI-LinUCB 25.9 1.24 6.9 8.8
CBSCFD 61.8 2.28 8.4 8.9
CBRAP 563.5 96.37 10.9 16.6
DBSLinUCB 518.8 1.5 7.3 9.1
Beauty PSI-LinUCB 83.6 3.5 21.7 26.8
CBSCFD 106.5 6.01 22.4 27.2
CBRAP 1566.0 350.0 41.7 67.0
DBSLinUCB 290.8 3.6 19.1 24.8
6.4 Batch size and rank trade-off
The batch size $B$ controls the frequency of PSI updates. Larger batches reduce training time but may slightly degrade quality. Figures 17(a) and 17(b) in Appendix show that training time decreases substantially with larger batches while hit rate remains stable. This enables flexible tuning: larger batches for latency-sensitive deployments, smaller batches for quality-critical scenarios.
(a) Amazon Health
(b) Amazon All Beauty
@@ 6.4 Batch size and rank trade-off
(a) Amazon Health
(b) Amazon All Beauty
Figure 7: Average Hit Rate for different $r$ for PSI-LinUCB.
@@ 7 Conclusion
We presented PSI-LinUCB, a scalable variant of LinUCB for large-scale contextual bandits. The method maintains a compact representation of the inverse regularized design matrix using a diagonal term and a low-rank correction, which allows efficient computation. Moreover, our method has an average complexity of updates of order $\mathcal{O}(dr)$ per candidate. A key direction for future work is to develop theoretical guarantees under direct approximations of the inverse regularized design matrix. To our knowledge, there are no regret guarantees for linear bandit algorithms in this setting. At the same time, as we show numerically, such algorithms might be preferable as compared to the classical sketching algorithms in particular applications.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A LinUCB with Batch Updates
Algorithm 5 LinUCB
1: for each arm $a\in\mathcal{A}$ set $A_{0,a}=\lambda I$
2:for t in {0,…,n-1} do
3:  $X_{t,a}=[]$, $R_{t,a}=[]$
4:  for all $(u,a,r)$ in $batch_{t}$ do
5:   $X_{t,a}.append\,[x_{t,a}]$
6:   $R_{t,a}.append\,[r]$
7:  end for
8:  for each arm $a$ in $batch_{t}$ do
9:   $A_{t+1,a}\leftarrow A_{t,a}+X_{t,a}^{\top}X_{t,a}$
10:   $b_{t+1,a}\leftarrow b_{t,a}+X_{t,a}^{\top}R_{t,a}$
11:   ${\theta}_{t+1,a}\leftarrow A_{t+1,a}^{-1}b_{t+1,a}$
12:  end for
13:end for
14:for each arm $a\in\mathcal{A}$ do
15:  Observe context $x_{t,a}$
16:  Compute $\text{UCB}_{t,a}={\theta}_{t,a}^{\top}{x}_{t,a}+\alpha\sqrt{{x}_{t,a}^{\top}{A}_{t,a}^{-1}{x}_{t,a}}$
17:end for
18:return $\arg\max_{a\in\mathcal{A}}\text{UCB}_{t,a}$
Appendix B Related Work and Baseline Selection
B.1 Scalable variants of LinUCB
Although LinUCB li2010contextual is a widely used algorithm in recommendation systems, the time and memory requirements increases with the dimension of the context $d$ and the number of items, since it is necessary to store and invert the matrix $A_{a}\in\mathbb{R}^{d\times d}$ for each item $a$. To address scalability constraints, several extensions and variants of the LinUCB algorithm have been proposed in the literature.
A widely used approach to accelerate LinUCB is to apply rank-1 updates of the ridge-regularized design matrix $A_{a}$ via the Sherman–Morrison identity Angioli2025Efficient, Ciucanu2022Implementing, Wang2022Dynamic, Yan2025CoCoB, Ozbay2024Comparative, zenati2022efficient. These updates reduce the per-round update complexity from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d^{2})$ while maintaining exact parameter estimates, but still require storing a full $d\times d$ matrix for each arm.
Another line of research aims to improve the efficiency of LinUCB through matrix sketching techniques. Early work kuzborskij2019efficient demonstrated that LinUCB can be efficiently implemented using the Frequent Directions (FD) sketching method, reducing the per-round update time from $\mathcal{O}(d^{2})$ to $\mathcal{O}(md)$, where $m$ is the sketch size and $d$ is the feature dimension. However, applying FD to contextual bandits has certain drawbacks. In particular, the FD sketching method violate the positive definite monotonicity design matrices $A_{t,a}$, which may affect stability and theoretical guarantees. To address this limitation, a Spectral Compensation Frequent Directions (SCFD) method and its adaptation (CBSCFD) for high-dimensional contextual bandit were proposed chen2020efficient, which preserves positive definiteness while maintaining the same $\mathcal{O}(md)$ computational and memory complexity. More recently, adaptive sketching techniques such as Dyadic Block Sketching wen2024matrix have been introduced to dynamically adjust the sketch size, ensuring a per-round update complexity of $\mathcal{O}(dl)$, where $l$ is the current sketch size. This adaptive strategy prevents excessive spectral loss and avoids linear regret when the spectrum of the design matrix decays slowly.
B.2 Baselines used in our experiments
To evaluate PSI-LinUCB, we select baselines that cover both exact implementations and implementations using covariance matrix approximation, which reduce either the effective rank of the matrix or the feature dimension.
As exact implementations, we report results for Batch LinUCB [li2010contextual], which matches batched training protocol provided in Algorithm˜5 and LinUCB Classic with Sherman–Morrison rank-one inverse updates [Angioli2025Efficient]. These baselines provide a useful quality reference and illustrate the memory and computational cost of exact updates in high dimensions.
To represent design-matrix compression via sketching, we use CBSCFD [chen2020efficient], which combines FD with an additional correction and maintains a low-rank sketch of the design matrix. CBSCFD is widely used as a robust sketching baseline for high-dimensional linear contextual bandits. We additionally include Dyadic Block Sketching (DBSLinUCB) [wen2024matrix] as a recent adaptive sketching approach that dynamically adjusts sketch size.
To cover feature-space compression, we include CBRAP [yu2017cbrap], which applies random projections before performing LinUCB-style updates. This baseline is conceptually different from sketching the design matrix and provides a typical alternative for dimensionality reduction.
Some methods are closely related but are not included as primary baselines. We do not include FD [kuzborskij2019efficient] in the main comparison because CBSCFD [chen2020efficient] combines FD with an additional correction and provides a stronger and more robust representative within the same sketching family. We also omit Diag-LinUCB [yi2023online] because it suggests to learn context representations (using non-linear transformations) followed by diagonal approximation and thus the comparison with original LinUCB is uninformative.
Appendix C Experiments
C.1 Hyperparameter Selection
Context dimensionality $d$
@@ C.1 Hyperparameter Selection
The optimal values of $r^{\prime}$ were selected using the scree plot, which displays singular values in descending order. The resulting context sizes are:
  • MovieLens 1M: $r^{\prime}=13$, $d=169$;
  • Magazine Subscriptions: $r^{\prime}=41$, $d=1681$;
  • Health & Personal Care: $r^{\prime}=37$, $d=1369$;
  • All Beauty: $r^{\prime}=51$, $d=2601$.
Regularization $\alpha$
@@
We used grid search to find the optimal exploration parameter $\alpha$. First, a broad grid $\alpha\in\{-10,-8,\dots,10\}$ was evaluated, followed by a refined search over $\alpha\in\{-1,-0.5,-0.3,-0.2,-0.1,-0.05,-0.03,-0.01,0.1\}$. The optimal $\alpha$ for LinUCB coincided with that for PSI-LinUCB.
Rank and sketch size
PSI-LinUCB rank, CBSCFD/CBRAP parameter $m$, DBSLinUCB block size $l_{0}$ (which is not fixed but grows exponentially during learning) and the error parameter $\varepsilon$, was explored over the grid $\{b\cdot 2^{i}\mid i\in\{1,\dots,8\},b\in\{2,3\}\}$.
Feature extraction
Given an interaction matrix $A\in\mathbb{R}^{m\times n}$, we apply truncated SVD: $A\approx U\Sigma V^{\top}$. Item features are obtained from $V^{\top}\in\mathbb{R}^{n\times r}$, and user features from $A\cdot V^{\top}\in\mathbb{R}^{m\times r}$. Features are extracted from warm-up data, cold users lacking valid context representations are excluded from evaluation. Note that cold-start users is not inherent to bandits, as soon as there are available contexts, the algorithm can be run, the restriction arises from our feature construction SVD-based pipeline and does not alter the fundamental bandit framework. Then we construct a context ${x}_{t,a}=\text{vec}(x_{u}{x}_{a}^{\top})\in\mathbb{R}^{d_{u}\times d_{a}}$, where $x_{u}$ and $x_{a}$ are user and item feature vectors. In our experiments, we set $d_{u}=d_{a}=r^{\prime}$, yielding $d=(r^{\prime})^{2}$.
Data processing
Training time measures initial warm-up training; evaluation time is the average prediction duration per user, averaged across test months. The interaction dataset is grouped by arms, with each batch corresponding to a fixed number of interactions for a single arm. This scheme is applied uniformly across all algorithms for fair comparison. The batch-to-rank ratio (see Remark 1) in PSI-LinUCB corresponds to the compression frequency factor of 2 in CBSCFD; however, batch size does not affect final model state or quality for CBSCFD and CBRAP, as these methods perform incremental per-sample updates.
(a) MovieLens 1M
(b) Magazine Subscriptions
Figure 8: Average Hit Rate for different $r$ for PSI-LinUCB.
Figure 9: Performance comparison across different context sizes on Amazon Health dataset.
(a) Health & Personal Care
(b) Magazine Subscriptions
(c) Beauty
@@
Figure 10: Algorithm scaling with number of arms on different datasets.
Figure 11: Online classification results on CIFAR-10.
Figure 12: Approximation error of $A^{-1}$ on CIFAR-10.
Table 6: Summary of datasets for online classification
Dataset #Samples #Features #Classes
MNIST 60000 784 10
CIFAR-10 50000 3072 10
(a) MovieLens 1M
(b) Amazon All Beauty
(c) Amazon Health
(d) Magazine Subscriptions
Figure 13: Average Hit Rate for different $m$ values for CBRAP.
(a) MovieLens 1M
(b) Magazine Subscriptions
Figure 14: Average Hit Rate for different $m$ for CBSCFD.
(a) MovieLens 1M
(b) Magazine Subscriptions
Figure 15: Average Hit Rate for different rank values for PSI-LinUCB.
(a) Amazon Health
(b) Magazine Subscriptions
Figure 16: Relative approximation error of $A^{-1}$ on real-world datasets.
C.2 Setting of online classification
The hyperparameters $\beta$ and $\lambda$ are selected via grid search over $\{10^{-4},10^{-3},\ldots,1\}$ and $\{2\times 10^{-4},2\times 10^{-3},\ldots,2\times 10^{4}\}$, respectively, following the experimental protocol of chen2020efficient. For our PSI-LinUCB algorithm, the rank parameter $r$ and for sketching-based methods (CBSCFD, CBRAP) the optimal sketch size $m$ are selected via cross-validation over a grid ranging from $10$ to $100$ with step $10$ and $\{200,300,400,500\}$ by minimizing the number of the average online mistakes on a validation run.
All experiments are repeated $20$ times and we report the average cumulative mistakes.
This dual applicability represents a practical advantage of our approach: the same algorithmic framework can be deployed both in the per-arm setting commonly used in real-world recommendation systems and in the shared model setting typical of experimental benchmarks. This allows for detailed theoretical analysis in controlled settings while maintaining direct applicability to production environments.
In addition to the baselines from chen2020efficient, we include a comparison with the Dyadic Block Sketching Linear (DBSL) algorithm wen2024matrix, which represents a more recent adaptive sketching approach.
Algorithm 6 PSI-LinUCB Rank-1 Update
0: context $x_{t}$, reward $r_{t}$, rank $r$
[…]
6:if $U_{t}.\text{shape}[1]<2r$ then
7:  $U_{t+1}\leftarrow[U_{t},\;\beta_{t+1}\bar{x}_{t+1}]$
8:  $V_{t+1}\leftarrow[V_{t},\;(I-V_{t}U_{t}^{\top})\bar{x}_{t+1}]$
9:end if
10:if $U_{t}.\text{shape}[1]=2r$ then
11:  if first time $U_{t}.\text{shape}[1]=2r$ then
12:   $\tilde{U}_{t+1}S_{t+1}\tilde{V}_{t+1}^{\top}=SVD(U_{t}V_{t}^{\top})$
13:  end if
14:  $\Delta D_{t+1}=\beta_{t+1}\tilde{x}_{t+1}\tilde{x}_{t+1}^{\top}(I-U_{t}V_{t}^{\top})$ {We do not form $\Delta D_{t+1}$ explicitly}
15:  $\tilde{U}_{t+1},S_{t+1},\tilde{V}_{t+1}=\textit{PSI}(\tilde{U}_{t},S_{t},\tilde{V}_{t},\Delta D)$
16:  $U_{t+1}\;\leftarrow\;\tilde{U}_{t+1}S_{t+1}$
17:  $V_{t+1}\;\leftarrow\;\tilde{V}_{t+1}$
18:end if
[…]
20:return $U_{t+1},V_{t+1},\theta_{t+1}$
(a) Amazon Health
(b) Amazon All Beauty
Figure 17: Batch/rank dependence on Amazon datasets.
Appendix D Proofs
D.1 Proof of Theorem˜1
Using (14), we obtain
$L_{t+1}^{-1}=(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})^{-1}L_{t}^{-1}\;.$ (35)
After applying the Sherman-Morrison-Woodbury formula
the recursive structure of the inverse update reads:
$L_{t}^{-1}=(I-\beta_{t}\bar{x}_{t}\bar{x}_{t}^{\top})L_{t-1}^{-1}\;.$ (37)
Unraveling this recursion from $t$ back to the initial condition gives
$L_{t+1}^{-1}$ $=(I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})L_{t}^{-1}$ (40)
This can be written as $L_{t+1}^{-1}=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1}$, where
$U_{t+1}$ $=\begin{bmatrix}U_{t}&\beta_{t+1}\,\bar{x}_{t+1}\end{bmatrix}$, (44)
This completes the proof.
D.2 Proof for Section˜2.2